Sean purchases office furniture from a wholesaler listed at $900, less discounts of 15% and 7%. He has overhead expenses of 16% of the cost and wants to have an operating profit of 30% of the cost.

a. Calculate the regular selling price of the office furniture.

b. After listing the furniture for one month, he marked it down by 19%. Calculate the profit or loss that he made at the reduced selling price.

c. What is the maximum rate of markdown that he can offer so that he breaks even on the sale?
Answer #1

Cost of furniture = 900(1-15%)(1-7%) = $711.45

Overheads = 16%*711.45 = $113.832

Total Cost = $825.282

a.Regular Selling price = 825.282(1+30%) = $1,072.87

b.Reduced Selling price = 1,072.87(1-19%) = $869.02

Profit = 869.02 - 825.28

= $43.74

c.Maximum mark down = (1,072.87-825.282)/1,072.87

= 23.08%